OpenStudy (anonymous):

How can you change 3x-y=1 into y=mx+b

OpenStudy (anonymous):

Move 3x to the other side using inverse operations. 3x-y=1 Subtract 3x from both sides -3x -3x -y=-3x+1

OpenStudy (anonymous):

Get it?

OpenStudy (anonymous):

Subtract 3x from both sides to get -y= -3x+1 Then, divide both sides by -1 to get rid of the negative on the y side You should get y=3x-1 M=3, b=-1

OpenStudy (anonymous):

Yea I get it thanks